# After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in RAG

### v1

@@ Abstract
Retrieval-augmented generation (RAG) systems face critical challenges in balancing internal (parametric) and external (retrieved) knowledge, especially when these sources conflict or are unreliable. To analyze these scenarios comprehensively, we construct the Trustworthiness Response Dataset (TRD) with 36,266 questions spanning four RAG settings. We reveal that existing approaches address isolated scenarios—prioritizing one knowledge source, naively merging both, or refusing answers—but lack a unified framework to handle different real-world conditions simultaneously. Therefore, we propose the BRIDGE framework, which dynamically determines a comprehensive response strategy of large language models (LLMs). BRIDGE leverages an adaptive weighting mechanism named soft bias to guide knowledge collection, followed by a Maximum Soft-bias Decision Tree to evaluate knowledge and select optimal response strategies (trust internal/external knowledge, or refuse). Experiments show BRIDGE outperforms baselines by 5–15% in accuracy while maintaining balanced performance across all scenarios. Our work provides an effective solution for LLMs’ trustworthy responses in real-world RAG applications.
@@ 1 Introduction
Recent advancements in retrieval-augmented generation (RAG) have enhanced the accuracy and trustworthiness of large language models (LLMs) in knowledge-intensive tasks, mitigating critical issues such as outdated model knowledge and hallucinations guu2020retrieval ; lewis2020retrieval . After retrieval, and before generating responses, LLMs often rely on two primary knowledge sources: context acquired through retrieval, referred to as external knowledge, and parametric knowledge stored within the LLM during training, termed internal knowledge. However, both knowledge sources exhibit limitations. External knowledge is prone to irrelevance, misinformation, or incompleteness due to corpus quality, retriever limitations, or query complexity chen2024benchmarking ; zou2024poisonedrag ; dai2024bias . For instance, in the NQ dataset kwiatkowski2019natural for open-domain question answering, wang2024astute employed Google Search as the retriever and searched the top 10 web pages for each question, finding that 34% of these pages were completely irrelevant to the answer. In the PopQA dataset mallen2022not , this percentage even reached 50%. Similarly, internal knowledge may be outdated or erroneous, as LLMs are constrained by training data noise and the rapid evolution of world knowledge weidinger2021ethical ; longpre2024pretrainer .
@@
In real-world RAG applications, the limitations of internal and external knowledge sources can lead to four distinct scenarios wang2024astute ; zhang2024evaluating (as illustrated in Figure 1). In the ideal case, LLMs should adopt scenario-specific strategies: (1) When both knowledge sources align and are correct, LLMs should be faithful to all knowledge to respond. (2) If external knowledge contains errors while internal knowledge remains reliable, LLMs should be faithful to internal knowledge. (3) Conversely, when internal knowledge is outdated or incorrect but retrieved context provides valid information, LLMs should anchor responses to external evidence. (4) When neither source contains sufficient or credible information, LLMs should refuse to answer.
Existing research addresses these scenarios separately. When facing scenario (2) or (3), some works advocate aligning LLM’s response with one of the knowledge sources (either internal or external knowledge) pan2023risk ; chen2024universal ; li2022large ; weller2022defending ; zhou2023context . Others attempt to integrate both sources shi2024trusting ; asai2023self ; xiang2024certifiably ; wang2024astute ; wei2025instructrag ; zhou2025trustrag but overlook scenario (4), failing to establish safeguards when both knowledge are invalid. A third line of work focuses exclusively on refusal mechanisms for exceeding the boundaries of knowledge sources cao2023learn ; song2025measuring , yet neglects knowledge integration strategies. In summary, these isolated approaches only use some specific strategies to handle one or two of these situations, failing to reveal the inadequacies of LLMs when faced with multiple RAG scenarios. Hence, there is also a lack of a synthesized framework capable of simultaneously addressing these diverse scenarios while ensuring LLMs’ trustworthiness in the RAG system.
To comprehensively evaluate LLMs’ response strategies under four RAG scenarios, we construct the Trustworthiness Response Dataset (TRD) containing 36,266 multiple-choice questions. Each question explicitly provides both external knowledge (simulating post-retrieval context) and internal knowledge (eliminating internal knowledge gaps between different LLMs). TRD comprises four RAG scenarios mirroring our proposed response strategies: (1) Faithful to All Knowledge (FA): We provide correct internal and external knowledge for the corresponding questions. (2) Faithful to Internal Knowledge (FI): We modify correct external knowledge with false facts to construct poisoned data. (3) Faithful to External Knowledge (FE): Standardizing data containing incorrect internal knowledge across LLMs from different families is complicated. Therefore, we use the latest Wikidata and Wikipedia facts (beyond the LLMs’ internal knowledge cutoff date) to construct questions and external knowledge. This ensures the models lack up-to-date internal knowledge, thereby forcing them to rely solely on external knowledge for answers. (4) Refuse to Answer (RA): While ensuring that internal knowledge is outdated, we introduce false facts to corrupt the correct external knowledge, rendering the questions unanswerable based on all knowledge sources. Our evaluation of TRD reveals that existing LLM families and RAG baselines fail to handle all scenarios effectively due to their lack of a unified understanding across these four cases.
@@
However, in this task, human behavioral patterns offer novel insights for addressing these scenarios. For instance, when confident in our internal knowledge, we perform limited external searches just for verification fu2006suboptimal . Conversely, for uncertain or time-sensitive questions, we conduct extensive retrieval and evaluate consistency of external knowledge rieh2002judgment . When no credible evidence exists, humans readily acknowledge ignorance and refuse to answer the question gigerenzer2010rationality —a critical behavior that current RAG systems struggle to emulate. Inspired by these observations, we propose Biased Retrieval anD Generation Evaluation (BRIDGE), a framework that enhances the trustworthiness of LLMs in RAG systems. Unlike prior work that either enforces strict alignment to single knowledge sources (hard bias) or treats all sources equally (no bias), BRIDGE introduces soft bias: an adaptive weighting mechanism where an Allocator module determines whether to prioritize broad external knowledge retrieval (for consistency evaluation) or deeper internal knowledge generation (for fact verification). Based on the TRD training set data, Allocator supports both in-context learning (ICL) and group relative policy optimization (GRPO) paradigms. Then, based on the weighted knowledge, BRIDGE computes matching scores and evaluates consistency through a Maximum Soft-bias Decision Tree—a bias-guided decision tree enabling fine-grained response strategy selection (e.g., trusting internal knowledge, following external evidence, or refusing to answer). Additionally, it incorporates a reflection strategy to rectify decision errors during inference. In summary, our contributions are as follows:
  • To comprehensively evaluate LLM response strategies in real-world RAG, we construct TRD and use it to reveal the limitations of prior isolated solutions.
  • We propose the BRIDGE, a unified approach introducing soft bias to balance internal and external knowledge dynamically, a Maximum Soft-bias Decision Tree, and a Reflection module—covering all scenarios. Furthermore, BRIDGE can be applied to both ICL and GRPO settings.
  • Experiments show that BRIDGE demonstrates superior and balanced performance on TRD and other trustworthiness benchmarks for single RAG scenarios.
2 Related Works
Retrieval-augmented Generation. The paradigm of RAG has evolved significantly since its inception, driven by the need to build LLMs on external knowledge while mitigating hallucinations. Early foundational work established the core framework of RAG guu2020retrieval ; lewis2020retrieval ; borgeaud2022improving . Subsequently, its retrieval module has been optimized through techniques such as query rewriting zheng2023take ; dai2023promptagator ; chan2024rqrag , passage reranking glass2022re2g ; yu2024rankrag , and adaptive efficient retrieval strategies asai2023self ; jiang2023active ; yao2023react . In the generation module, recent efforts have focused on improving generation efficiency via prompt compression jiang2024piperag ; xu2024recomp ; cheng2024xrag and decoding optimization merth2024superposition ; jin2024ragcache ; wang2025speculative . These approaches primarily enhance individual RAG components in non-adversarial settings, whereas our research focuses on adversarial conditions (internal and external knowledge containing conflicts or noise).
Trustworthiness of LLMs in RAG. Ensuring the trustworthiness of LLMs in the RAG system has become a critical research area, particularly in adversarial environments. For scenarios where internal and external knowledge sources conflict, some studies align model responses with internal knowledge pan2023risk ; xu2023earth ; hong2024so ; weller2022defending , while others prioritize external knowledge li2022large ; gekhman2023trueteacher ; zhou2023context . Beyond these binary choices, recent efforts to improve the trustworthiness of LLM responses attempt to merge conflict sources through contrastive decoding jin2024tug ; shi2024trusting or integrating consistent portions of internal and external knowledge xiang2024certifiably ; wang2024astute ; wei2025instructrag ; zhou2025trustrag . However, these methods rely on carefully designed prompt templates to assess knowledge consistency. This matching strategy is sensitive to textual variations and lacks mechanisms for handling unanswerable questions. Previous work on enhancing LLM refusal capabilities in RAG systems relies on building expert knowledge bases cao2023learn or fine-tuning generative models song2025measuring . However, these methods often fail to integrate multiple knowledge sources and lack plug-and-play compatibility with black-box model-based RAG systems. Collectively, these works are orthogonal to each other, typically targeting only one or two scenarios. In contrast, BRIDGE provides a unified framework that systematically addresses four RAG scenarios through soft bias allocation and a scored decision tree.
@@ 3 TRD Construction
In this section, we explain the construction process of TRD. We formalize our task scenario as follows: given a natural language question $q$, the system retrieves external knowledge $K_{ext}$ and generates explicit internal knowledge $K_{int}$ by answering $q$ without accessing $K_{ext}$. Each question in TRD is associated with four options: one correct option, three incorrect options. One of the options is "I don’t know".. The dataset is divided into a training set (29,012 questions), a validation set (3,627 questions), and a test set (3,627 questions). TRD simulates four RAG scenarios displayed in Figure 1. Below, we detail the construction process.
@@ 3.1 Data Preparation
We utilize two primary datasets to help construct multiple-choice questions in TRD. (1) NQ: NQ is an open-domain question answering dataset released by Google in 2019 kwiatkowski2019natural . The accessible dataset contains 315,203 question-answer pairs (QA pairs) and provides a long answer (collected from Wikipedia by annotators) and a short answer (one or more entities) for each question. We use NQ to generate questions for the FA and FI categories. (2) TAQA: A dataset of 20,148 time-sensitive QA pairs derived from Wikipedia tables, where the answers to the questions change over time zhao2024set . We use TAQA to construct part questions for the FE and RA categories.
@@ 3.2 Question Generation
Faithful to All Knowledge. We select QA pairs from the NQ training set that contain long answers and short answers that are not timestamps. The short answer serves as the correct option. Using a popularity-based entity substitution method longpre2021entity , we generate two incorrect answers that have the same entity type as the short answer from Wikidata. The long answer is treated as external knowledge, while GPT-4-turbo generates internal knowledge based on the question and the short answer. We believe that explicitly providing internal knowledge helps to eliminate the parametric knowledge gap among different family LLMs during evaluation.
Faithful to Internal Knowledge. Following the same procedure as above, we select questions from the NQ validation set with long answers and non-temporal short answers, and generate two incorrect options of the same entity type using the popularity-based entity substitution method longpre2021entity . However, in this setting, we replace the correct entity in the long answer with one of the incorrect entities, resulting in misleading external knowledge. The internal knowledge, generated by GPT-4-turbo from the original short answer, remains correct.
Faithful to External Knowledge. To simulate scenarios where LLMs must rely solely on external knowledge due to outdated or absent internal knowledge, we define two distinct question sub-types: (1) Evolution: This category targets factual updates for existing facts/events (e.g., Who won the latest Nobel Prize in Literature?). LLMs trained with a cutoff date (February 2023 in this paper) cannot track such updates. The TAQA dataset focuses on this type of question. For each question, we query its latest answer in Wikipedia (February 1, 2025 dump). We retain only questions with valid Wikipedia pages. Then, using prompt templates in ConflictBank su2024conflictbank , we we use GPT-4-turbo to generate a Wikipedia-style evidence paragraph from the correct answer (post-cutoff date) as external knowledge. For the same question, we generate a parallel evidence paragraph describing TAQA’s 2021 answer (pre-cutoff date) using the same templates, simulating outdated internal knowledge. The correct answer is the entity with the post-cutoff date. We use the 2020 and 2021 answers in TAQA as incorrect options. (2) Perpetuation: This category addresses entirely new post-cutoff entities/events (e.g., What is the CPU of iPhone 16?). For many LLMs whose training data ends in 2023, it is unknown. First, we execute SPARQL [^1] https://www.wikidata.org/wiki/Wikidata:SPARQL_tutorial queries on Wikidata to extract timestamped triples where the object entity was created after February 2023 (e.g., <iPhone16, CPU, A18>). For these triples, we retrieve Wikipedia paragraphs (February 1, 2025 dump) containing both the subject and object entities as external knowledge; triples without corresponding paragraphs are filtered out. Internal knowledge is explicitly set to "None" to eliminate parametric knowledge interference from different LLMs. We select half of the data as FE data, and the other half is prepared to generate RA data.
@@
Refuse to Answer. We replace the correct answer in external knowledge with an outdated distractor entity (e.g., altering A18 to A17 in the iPhone16 example). This ensures no credible evidence exists across all knowledge sources, asking LLMs to abstain.
@@ 3.3 Question Validation
We use GPT-4-turbo and Llama3 70B to validate questions, ensuring:(1) For FA, both models could select correct answers using either knowledge source. (2) For FI, models must choose the correct answer using internal knowledge but a consistent incorrect answer with external knowledge. (3) For FE, models must answer correctly with external knowledge but fail with internal knowledge. (4) For RA, both models select wrong answers using either knowledge source.
After validation, we balance the dataset, limiting FA to 10% of the total (since it is not the primary focus of our research) and allocating 30% each to the other three categories. Furthermore, the construction of time-related questions is automated, allowing them to be updated based on changes to Wikipedia and Wikidata, thus meeting the future evaluation needs of LLMs with evolving internal knowledge. For TRD construction details and statistics, please see Appendix A.
@@ 4 Biased Retrieval and Generation Evaluation
As illustrated in Figure 2, BRIDGE operates through two sequential phases: (1) Bias-Guided Knowledge Collection: This phase constructs an Allocator that computes retrieval dependency probability $r_{p}$ and generation dependency probability $g_{p}$ to determine the reliance degree on external and internal knowledge (constrained by $r_{p}+g_{p}=100\%$). We implement two Allocator paradigms: a reinforcement learning-based approach using GRPO and ICL method (§ 4.1.1). Subsequently, the system generates $n$ sub-queries capturing key information of $q$, and allocating the number of sub-queries based on $r_{p}$ and $g_{p}$. For these sub-queries, we yield generated knowledge $K_{gen}$ from Multi-query Generator and retrieved knowledge $K_{ret}$ from Multi-query Retriever (§ 4.1.2). (2) Bias-Guided Knowledge Evaluation: This phase employs a BGE-m3 chen2024bge based encoder as a Scorer to compute multi-granularity matching scores between knowledge sources (§ 4.2.1). These scores feed into a Maximum Soft-bias Decision Tree that determines the optimal response strategy under bias weights and thresholds. Furthermore, a Reflection module will trigger re-execution when the decision tree detects potential decision errors (§ 4.2.2).
Figure 2: Overview of BRIDGE.

### v2

# After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in Retrieval-Augmented Generation

@@ Abstract
Retrieval-augmented generation (RAG) is a promising paradigm, yet its trustworthiness remains a critical concern. A major vulnerability arises prior to generation: models often fail to balance parametric (internal) and retrieved (external) knowledge, particularly when the two sources conflict or are unreliable. To analyze these scenarios comprehensively, we construct the Trustworthiness Response Dataset (TRD) with 36,266 questions spanning four RAG settings. We reveal that existing approaches address isolated scenarios—prioritizing one knowledge source, naively merging both, or refusing answers—but lack a unified framework to handle different real-world conditions simultaneously. Therefore, we propose BRIDGE, a framework that dynamically determines comprehensive response strategies for large language models (LLMs). BRIDGE leverages an adaptive weighting mechanism to guide knowledge collection, followed by a decision tree to evaluate evidence and select the optimal response strategy (e.g., trusting internal/external knowledge, or refusal). Experiments show BRIDGE outperforms baselines by 5–15% in accuracy while maintaining balanced performance across all scenarios. Our work provides an effective solution for producing trustworthy LLM responses in RAG applications.  [^1] The code and benchmark is publicly available at https://github.com/Kangkang625/BRIDGE.
After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in Retrieval-Augmented Generation
Xinbang Dai†∗   Huikang Hu†∗   Yuncheng Hua   Jiaqi Li   Nan Hu Yongrui ChenRihui JinYuxin ZhangYuyang Zhang Xiaoguang LiLifeng ShangGuilin Qi Southeast University  University of New South Wales  Noah’s Ark Lab {xbdai, huikanghu}@seu.edu.cn
@@ 1 Introduction
Recent advancements in retrieval-augmented generation (RAG) have enhanced the accuracy of large language models (LLMs) in knowledge-intensive tasks Guu et al. (2020); Lewis et al. (2020). After retrieval and before response generation, LLMs often rely on two primary knowledge sources: context acquired through retrieval, referred to as external knowledge, and parametric knowledge stored within the LLM during training, termed internal knowledge. However, both sources have limitations that can undermine the trustworthiness of RAG systems. Specifically, the retrieval process may introduce irrelevant or poisoned content Chen et al. (2024b); Zou et al. (2024), whereas internal knowledge may be outdated or erroneous due to temporal limitations or noise in the training data Weidinger et al. (2021); Longpre et al. (2024).
@@
In real-world RAG applications, limitations in internal and external knowledge sources can lead to four distinct scenarios Wang et al. (2024); Zhang et al. (2024) (as illustrated in Figure 1). Ideally, LLMs should adopt scenario-specific strategies: (1) When both knowledge sources are consistent and correct, LLMs should base their responses on all available knowledge. (2) If external knowledge contains errors while internal knowledge is reliable, LLMs should rely on their internal knowledge. (3) Conversely, when internal knowledge is outdated or incorrect but the retrieved context is valid, LLMs should anchor responses to external evidence. (4) When neither source contains sufficient or credible information, LLMs should refuse to answer.
Unfortunately, existing research typically addresses these scenarios in isolation. Some studies enforce alignment of LLM responses with a single source (either internal or external) Pan et al. (2023); Chen et al. (2024c); Li et al. (2023); Weller et al. (2024); Zhou et al. (2023), whereas others integrate both sources but overlook the need for refusal mechanisms (Scenario 4) Shi et al. (2024); Asai et al. (2023); Xiang et al. (2024); Wang et al. (2024); Wei et al. (2025); Zhou et al. (2025). Conversely, safety-focused works discuss refusal but neglect effective knowledge integration Cao (2024); Song et al. (2025). As a result, existing frameworks do not simultaneously consider all four scenarios required for trustworthy RAG.
To systematically evaluate LLM response strategies across these four scenarios, we introduce the Trustworthiness Response Dataset (TRD), comprising 36,266 multiple-choice questions. TRD covers four categories, each aligned with a desired response strategy: (1) Faithful to All Knowledge (FA): both internal and external knowledge support the correct answer, so models should base responses on all available knowledge; (2) Faithful to Internal Knowledge (FI): external context is deliberately poisoned while internal knowledge remains reliable, so models should rely on internal knowledge; (3) Faithful to External Knowledge (FE): questions require up-to-date facts beyond the model’s pre-training cutoff dates, so internal knowledge is insufficient and models must rely on external evidence; and (4) Refuse to Answer (RA): both internal and external knowledge are unreliable, so models should refuse to answer.
Our evaluation on TRD shows that when the model is simultaneously exposed to two potentially imperfect knowledge sources, it may adopt unreasonable response strategies. Moreover, existing LLM families and RAG baselines exhibit refusal behavior in fewer than 10% of cases (as shown in Table 1). Human behavioral patterns offer inspiration: humans verify known facts with minimal retrieval Fu and Gray (2006), deeply evaluate external sources for uncertain or time-sensitive questions Rieh (2002), and readily admit ignorance when evidence is lacking Gigerenzer (2010).
@@
Therefore, we propose Biased Retrieval anD Generation Evaluation (BRIDGE), a framework that enhances the trustworthiness of LLMs in RAG systems. Unlike prior work that either enforces strict alignment with a single knowledge source (hard bias) or treats all sources equally (no bias), BRIDGE introduces soft bias: an adaptive weighting mechanism where an Allocator module determines whether to conduct deeper internal knowledge generation (for fact verification) or prioritize broad external knowledge retrieval (for consistency evaluation). Based on the TRD training data, Allocator supports both in-context learning (ICL) and group relative policy optimization (GRPO) paradigms. Subsequently, based on the weighted knowledge sources, BRIDGE computes matching scores and evaluates consistency through a Maximum Soft-bias Decision Tree—an interpretable decision tree enabling fine-grained response strategy selection (e.g., trusting internal knowledge, following external evidence, or refusing to answer). Additionally, it incorporates a reflection strategy to rectify decision errors during inference. In summary, our contributions are as follows:
  • We construct TRD, a comprehensive benchmark designed to evaluate LLM response strategies across diverse RAG scenarios.
  • We propose BRIDGE, a unified framework that addresses diverse RAG scenarios.
  • Experiments demonstrate that BRIDGE achieves superior and more balanced performance compared to existing methods, not only on TRD but also on established benchmarks targeting individual RAG scenarios.
@@ 2 TRD Construction
This section details the construction process of TRD. We formalize our task scenario as follows: given a natural language question $q$, the system retrieves external knowledge $K_{ext}$ and generates explicit internal knowledge $K_{int}$ by answering $q$ without accessing $K_{ext}$. Each question in TRD is associated with four options: one correct option and three incorrect options. One option is "I don’t know," allowing the model to abstain from answering. The dataset comprises 29,012 questions for training, 3,627 for validation, and 3,627 for testing.
@@ 2.1 Data Preparation
We use two datasets to build TRD’s multiple-choice questions. (1) NQ Kwiatkowski et al. (2019): An open-domain QA dataset with 315,203 QA pairs, each paired with a long answer (Wikipedia passage) and one or more short answer entities. We use NQ to generate questions for the FA and FI categories. (2) TAQA Zhao et al. (2024): A dataset of 20,148 time-sensitive QA pairs from Wikipedia tables, where answers change over time. We use TAQA to construct part of the FE and RA questions.
@@ 2.2 Question Generation
Faithful to All Knowledge (FA). From the NQ training set, we designate the short answer as the correct option for each question. Using a popularity-based entity substitution method Longpre et al. (2021), we generate two incorrect answers from Wikidata that share the same entity type as the short answer. For each question, the corresponding long answer serves as the external knowledge $K_{ext}$. To obtain the internal knowledge for all question types (FA, FI, FE, RA), we prompt three backbone models (GPT-3.5-turbo, Qwen-72B, and Llama-3-8B-Instruct) to answer each question directly. We regard the generated responses as explicit representations of the models’ internal knowledge.
Faithful to Internal Knowledge (FI). Questions of type FI are derived from the NQ validation set. The generation procedure for answer options follows that of FA. To construct misleading external knowledge, we replace the correct entity in the long answer with one of the incorrect entities from the distractor options, resulting in a corrupted $K_{ext}$ that introduces factual conflict while preserving textual plausibility.
Faithful to External Knowledge (FE). To simulate scenarios where LLMs must rely solely on external knowledge due to outdated or absent internal knowledge, we define two distinct types:
(1) Evolution: This category covers questions that require updated information about existing facts and events (e.g., Who won the latest Nobel Prize in Literature?). LLMs with a training data cutoff (in our paper, February 2023) are unable to track such updates. Following the methodology described in the TAQA dataset paper, we update the short answers and their corresponding contexts using a Wikipedia dump dated February 1, 2025. The context extracted from Wikipedia serves as external knowledge $K_{ext}$, with the updated short answers treated as correct options. The original answers from the TAQA dataset (dated 2020 and 2021) are retained as incorrect distractors.
(2) Perpetuation: This category addresses new entities or events that have emerged after the LLM’s knowledge cutoff (e.g., What is the CPU of iPhone 16?). For many LLMs whose training data ends in 2023, such information remains unknown. First, we execute SPARQL [^2] https://www.wikidata.org/wiki/Wikidata:SPARQL_tutorial queries on Wikidata to extract timestamped triples created after February 2023 (e.g., <iPhone 16, hasCPU, A18>). For each of these triples, we retrieve textual paragraphs from Wikipedia that contain both the subject and object entities as external knowledge. We then mask the object entity of the triple and prompt GPT-4-turbo to generate a question for which the masked entity serves as the correct answer. In addition, we retrieve two incorrect distractor options using the substitution method Longpre et al. (2021). It is worth noting that when relying on internal knowledge, the model is unaware of these new facts; consequently, its responses typically consist of hallucinated content or explicit admissions of ignorance. Finally, we split the resulting dataset evenly: one half is used as FE data, and the other half is used to construct subsequent RA data.
@@
Refuse to Answer (RA). We corrupt the external knowledge text by replacing the correct answer with a distractor entity (e.g., changing A18 to A17 in the iPhone 16 example). This ensures that no credible evidence exists in any knowledge source, creating a scenario in which the model should ideally refuse to answer. Finally, we set the correct answer to "I don’t know."
Figure 2: Overview of BRIDGE.
@@ 2.3 Question Validation
To validate the questions in our TRD benchmark, we use the backbone LLMs employed in this study for verification and filtering. This validation process adheres to a strict set of criteria for each question category. (1) For FA, a question is retained only if the models can answer it correctly when given only the internal knowledge source or only the external knowledge source. (2) For FI, a question is retained only if the models answer it correctly using internal knowledge alone but, when provided with erroneous external information, are misled into selecting the target incorrect answer. (3) For FE, a question is retained if the models fail to produce the correct answer when relying exclusively on internal knowledge. (4) For RA, a question is retained if the models cannot derive the correct answer from either knowledge source. For additional details on TRD construction and statistics, please refer to Appendix A.
@@ 3 Biased Retrieval and Generation Evaluation
As illustrated in Figure 2, BRIDGE processes the user question $q$ in two stages: (1) Bias-Guided Knowledge Collection: The Allocator first estimates how much the question $q$ relies on internal versus external knowledge, producing the generation and retrieval dependency probabilities ($g_{p}$ and $r_{p}$, where $r_{p}+g_{p}=100\%$) (§ 3.1.1). The system then generates $n$ sub-queries that capture key information from $q$, and allocates them to the Multi-query Generator and the Multi-query Retriever according to $g_{p}$ and $r_{p}$, yielding generated knowledge $K_{gen}$ and retrieved knowledge $K_{ret}$ (§ 3.1.2). (2) Bias-Guided Knowledge Evaluation: The framework evaluates four knowledge sources: self-generated internal knowledge $K_{int}$ , initially retrieved external knowledge $K_{ext}$, and the previously obtained $K_{gen}$ and $K_{ret}$. We employ the BGE-m3 encoder Chen et al. (2024a) to compute multi-granularity similarity scores across knowledge types (§3.2.1). Then, a Maximum Soft-bias Decision Tree uses these scores to determine the optimal response strategy. Moreover, if potential errors are detected, a reflection module triggers re-execution for refinement (§3.2.2).